OpenStudy (anonymous):

.Brandon needs $480 to buy a TV and stereo system for his room. He received $60 in cash for birthday presents. He plans to save $30 per week from his part-time job. To find how many weeks w it will take to have $480, solve 60 + 30w = 480.Type your answer in the blank to the left.

OpenStudy (ivettef365):

just solve for x

OpenStudy (anonymous):

I'm doing this..the seventh grader way! >:D This is very easy. All you do is the opposite, i think. I'm not a smart person. Subtract 480 from 60 and when you get that answer divide it by 30. I'm pretty sure It'd give you the right answer but I'm sorry if you get it wrong.

OpenStudy (ivettef365):

that should work

OpenStudy (anonymous):

haha i know its simple but my teacher said he threw some thing else in here to trick. or he might just be tricking the class with that statement. i just needed to see if there was some thing i didnt see. thanks! :)

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (texaschic101):

60 + 30w = 480 30w = 480 - 60 30w = 420 w = 420/30 w = 14 weeks <--answer